# Advances in Robust Federated Learning: Heterogeneity Considerations

### v1

@@ Abstract
In the field of heterogeneous federated learning (FL), the key challenge is to efficiently and collaboratively train models across multiple clients with different data distributions, model structures, task objectives, computational capabilities, and communication resources. This diversity leads to significant heterogeneity, which increases the complexity of model training. In this paper, we first outline the basic concepts of heterogeneous federated learning and summarize the research challenges in federated learning in terms of five aspects: data, model, task, device and communication. In addition, we explore how existing state-of-the-art approaches cope with the heterogeneity of federated learning, and categorize and review these approaches at three different levels: data-level, model-level, and architecture-level. Subsequently, the paper extensively discusses privacy-preserving strategies in heterogeneous federated learning environments. Finally, the paper discusses current open issues and directions for future research, aiming to promote the further development of heterogeneous federated learning.
@@
Federated Learning, heterogeneity, robust, privacy-preserving
@@ I Introduction
With the popularization of mobile devices and 5G networks, a large amount of data is generated on edge devices every day, greatly promoting the development of artificial intelligence [1]. Traditional machine learning (ML) methods require centralized data storage and processing for model training, which poses significant challenge to data transmission and storage. Moreover, with the frequent occurrence of data leakage, the issue of privacy protection attracts more and more attention. In this context, institutions and enterprises are going to seek a new ML paradigm, which can utilize distributed data resources for training while protecting data privacy and security.
@@
In 2016, Google researchers proposed the concept of federated learning (FL) [2]. The core idea about FL is to train and update models locally on data owners (i.e. clients) and aggregates model updates on a third party (i.e. server), without the need to transfer the raw data to a central server. By this way, FL can significantly reduce data transmission in the network, thus reducing the possibility of privacy leakage and the requirement for centralized computing and storage resources. FL not only presents the respect and protection for privacy, but also demonstrates how to maximize the utility of data while ensuring privacy in the era of big data. At present, FL is becoming an effective way for balancing the utility and privacy of data, bringing new development directions and opportunities to ML. FL has demonstrated its broad application potential in fields such as intelligent healthcare [3] finance [4], the Internet of Things [5], and smart cities [6].
@@
However, in the application scenarios of typical FL, there are multiple heterogeneity issues among clients, including data heterogeneity, model heterogeneity, task heterogeneity, communication heterogeneity, and device heterogeneity, which brings challenges to collaboratively model training. The presence of heterogeneity not only significantly damages the generalization ability and prediction accuracy of trained model, but also to some extent restrict the application of FL. Specifically, FL often counters the following heterogeneity: Data heterogeneity: The clients often far away from each other geographically and lack communications, and each client may collect data from different data source, causing skews in the distributions of data collected by clients [7]. Moreover, the method of sampling and processing data and the amount of collected data may also vary across different clients, resulting in quality skew and quantity skew. Model heterogeneity: Due to device limitations, task types and other reasons, the clients have to customize their own model structures and jointly train models in a federated manner, which is model structure heterogeneity [8]. Moreover, the adopted optimizer of different clients may also be different. Task heterogeneity: Sometimes different clients need to train models for different tasks. For example, some clients may need to train models for image classification, while some other clients may need to train models for semantic segmentation.[9]. Communication heterogeneity: In typical FL, each round of FL requires transmission of model parameters or gradients between the server and the clients. But in reality, the participants of FL are usually under different network environments, and communication bandwidth and budgets of some participants are limited, which is called communication heterogeneity [10]. Device heterogeneity: The typical FL algorithms usually require local training on the clients. But due to limited computing power, storage resources, energy supply, etc., the clients may update local models inconsistently and thus lower the efficiency of the whole system [11].
@@
Research addressing heterogeneous FL is currently generating a lot of buzz, and we investigate a number of frequently cited or recent surveys related to heterogeneity and privacy issues. Although they have all provided detailed descriptions and abundant related researches for their respective focused topics, they still overlook some less mentioned types of issues and techniques. In this paper, we first categorize several types of heterogeneity in detail, and explore heterogeneity challenges may encountered while applying FL for practical environments, as well as solutions. For each type of heterogeneity, we not only present detailed and formal description, but also conduct extensive survey for comprehensively summarizing the state-of-the-art solutions. Moreover, given that privacy protection is one of the core demands of FL, we also focus on privacy protection problems and solutions while applying FL in scenarios with heterogeneity. We discuss the application of existing privacy protection techniques in FL, delving into the trade-off between privacy and utility, and introduce the advanced researches in relevant fields. We list the differences between our survey and the above other surveys in Table I.
@@
Survey [16] [17] [18] [19] [11] [20] [21] [22] [23] Ours
Data Heterogeneity Fig. 1 Distribution Skew $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Label Skew $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$
Feature Skew $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Quality Skew $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$
Quantity Skew $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$
Model Heterogeneity Fig. 2 Structure Heterogeneity $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$
Optimizer Heterogeneity $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$
Task Heterogeneity Fig. 3 $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
Communication Heterogeneity $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Device Heterogeneity $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Privacy Mechanism Fig. 4 Differential Privacy $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Homomorphic encryption $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Model Watermark $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$
Blockchain $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
Thus, the main contributions of this work are threefold:
@@
    We systematically categorize and analyze key components in heterogeneous FL, including data, model, task, communication, and device heterogeneity. For each type of heterogeneity, we discuss in detail the algorithmic challenges it raises.
@@
    We categorize existing state-of-the-art approaches into data-level, model-level, and architecture-level categories based on the level of processing, and discuss these classifications in depth.
@@
    We summarize the current privacy-preserving strategies in heterogeneous FL and explore possible future research directions with a view to promoting a safer and more efficient development of the field.
@@ II Classification and Challenges in Heterogeneous Federated Learning
In this section, we first provide a typical federation learning formulation to illustrate the typical federation learning process through an example. Federated learning is a machine learning setup in which multiple clients collaboratively train a model while protecting data privacy [24]. Each participant trains the model on local data and then sends only updates to the model (e.g., weights or gradients) to a central server. The server aggregates these updates and sends the aggregated results back to the participants to update their local models in the next iteration. In this way, federated training produces a global model that performs well on the entire dataset without sharing any raw data. In a classic FL framework [2], assuming that there are $K$ clients participating in the training, for $k$-th client $C_{k}$ has a private dataset $D_{k}=\left\{(x_{i}^{k},y_{i}^{k})\right\}_{i=1}^{N_{k}}$ that satisfies $|x^{k}|=N_{k}$ and $N=\sum_{k=1}^{K}N_{k}$. The local distribution of the data of client $k$ is denoted as $P_{k}(x,y)$, where $x$ and $y$ denote the features of the data samples and the labels. Specifically, FL can be formally defined as the following optimization problem:
@@
where $\theta$ is the global model parameter. $F_{k}(\theta)$ is the local loss function on client $k$, which reflects the performance of the model on that client. $p_{k}$ is the client weight, the value of which depends on the amount of data on the client and fulfills $p_{k}=\frac{N_{k}}{N}$ and $\sum_{k=1}^{K}p_{k}=1$. Therefore, the training process of FL can be represented as the following steps:
@@
    $\mathrm{Model\ parameters:}\theta\leftarrow\sum_{k=1}^{N}p_{k}\theta_{k},$ (2)
    $\mathrm{\textbf{or}\ Model\ gradient:}\theta\leftarrow\theta+\sum_{k=1}^{N}p_{k}\Delta\theta_{k}.$
@@
Although FL has been widely used in various domains [1], real-world FL usually faces heterogeneity in data distribution, model architecture, task requirements, network environments, and hardware devices among participant clients [16]. These heterogeneities pose a number of challenges, including how to effectively deal with differences in data distribution across clients, how to adapt to the computational and storage capabilities of different clients while maintaining model performance, how to balance knowledge sharing with task specificity in multitasking scenarios, how to optimize communication strategies to cope with varying bandwidth and latency conditions, and how to adapt to the energy and resource constraints of different devices. We next analyze the challenges posed by these heterogeneities, the resolution of which is essential to achieve efficient, scalable, and robust FL systems.
@@ II-A Data Heterogeneity
Data heterogeneity refers to the inconsistency of data distribution among clients in FL, which does not obey the same sampling, i.e., the data are not independently and identically distributed (Non-IID) [7, 23]. To explore data heterogeneity, we categorize data heterogeneity from the perspective of client data distribution. Specifically, we distinguish different categories of Non-IID data by five different skew patterns: distribution skew, label skew, feature skew, quality skew, and quantity skew, as shown in Fig. 1. Numerous studies[25, 26] have found that the client’s local optimization objective is inconsistent with the global optimization objective due to differences in the client’s local data distribution. As a result, data heterogeneity may cause the local model to converge in different directions to reach the local optimum rather than the global optimum, which reduces the FL performance and leads to poor global convergence of the model, and these problems deteriorate the performance of the global model for each participant, and may prevent participants from engaging in FL [27].
@@ II-A1 Distribution Skew
In FL, data distribution skew refers to the skew in the distribution of data held by different clients, i.e., the distribution of features or joint distribution of data from different clients is inconsistent with the overall data distribution [28]. In this case, the data no longer obeys the assumption of independent identically distributed (IID). As shown in Fig. 1a, there is $P_{i}(x,y)\neq P_{j}(x,y)$ due to the difference in the data distribution of clients $i$ and $j$.
A typical example is data obeying a Dirichlet distribution [29]. The Dirichlet distribution is a multivariate probability distribution that is often used to describe the distribution of the proportions of each category in a multicategorization problem. In the context of FL, the Dirichlet distribution can be used to model differences in the distribution of data labels on different clients. Where the parameter $\beta$ determines the degree of distributional difference of the data. As shown in the scatterplot in Fig. 1a, which illustrates the visualization of the difference in the distribution of data across 20 clients for different values of $\beta$.
@@ II-A2 Label Skew
Data label skew is a common Non-IID case [30], which refers to the skew in the distribution of data labels held by different clients. This label distribution skew means that even if the feature distributions are shared, the label distributions of different clients may be different [31], i.e., for clients $i$ and $j$, their label distributions $P_{i}(y)\neq P_{j}(y)$. However, for each label, the feature distributions of the data samples are similar regardless of which client they belong to, i.e., $P_{i}(x|y)=P_{j}(x|y)$. For example, in the data label skewing scenario shown in Fig. 1b, different users contain data from the set of handwritten digits, but each client may mainly contain a certain number of digits. The scatterplot shows a visualization of label skew for 20 clients. This data label skew can lead to a large difference in the gradient of the model uploaded back from different clients during federated training, thus affecting the convergence and generalization performance of the model.

### v2

# Advances in Robust Federated Learning: A Survey with Heterogeneity Considerations

@@ Abstract
In the field of heterogeneous federated learning (FL), the key challenge is to efficiently and collaboratively train models across multiple clients with different data distributions, model structures, task objectives, computational capabilities, and communication resources. This diversity leads to significant heterogeneity, which increases the complexity of model training. In this paper, we first outline the basic concepts of heterogeneous FL and summarize the research challenges in FL in terms of five aspects: data, model, task, device and communication. In addition, we explore how existing state-of-the-art approaches cope with the heterogeneity of FL, and categorize and review these approaches at three different levels: data-level, model-level, and architecture-level. Subsequently, the paper extensively discusses privacy-preserving strategies in heterogeneous FL environments. Finally, the paper discusses current open issues and directions for future research, aiming to promote the further development of heterogeneous FL.
@@
Federated Learning, heterogeneity, robust, privacy-preserving, survey
@@ I Introduction
With the popularization of mobile devices and 5G networks, a large amount of data is generated on edge devices every day[1]. Traditional machine learning (ML) methods require centralized data storage and processing for model training, which poses significant challenge to data transmission and storage. Moreover, with the frequent occurrence of data leakage, the issue of privacy protection attracts more and more attention. Thus, institutions and enterprises are looking for a new ML paradigm, which can utilize distributed data resources for training while protecting data privacy and security.
@@
In 2016, Google researchers proposed the concept of federated learning (FL) [2]. The core idea about FL is to train and update models locally on data owners (i.e. clients) and aggregates model updates on a third party (i.e. server), without the need to transfer the raw data to a central server. By this way, FL can significantly reduce data transmission in the network, thus reducing the possibility of privacy leakage and the requirement for centralized computing and storage resources. FL not only presents the respect and protection for privacy, but also demonstrates how to maximize the utility of data while ensuring privacy in the era of big data. At present, FL is becoming an effective way for balancing the utility and privacy of data, bringing new development directions and opportunities to ML. FL has demonstrated its broad application potential in fields such as intelligent healthcare [3], finance [4], the Internet of Things [5], and smart cities [6].
@@
However, in the application scenarios of typical FL, there are multiple heterogeneity issues among clients, which brings challenges to collaboratively model training. The presence of heterogeneity not only significantly damages the generalization ability and prediction accuracy of trained model, but also to some extent restrict the application of FL. Specifically, FL often counters the following heterogeneity: Data heterogeneity includes several types of skews in the data collected by clients. Before FL, each client may collect data from different data source, which may result in distribution skew, label skew and feature skew [7, 8]. Moreover, the method of sampling and processing data and the amount of collected data may also vary across different clients, resulting in quality skew and quantity skew. Model heterogeneity includes model structure heterogeneity and model optimizer heterogeneity. Model structure heterogeneity means that the clients have to customize their own model structures and jointly train models in a federated manner, due to device limitations, task types and other reasons [9]. Model optimizer heterogeneity means that the adopted optimizer of different clients may also be different. Task heterogeneity means that different clients need to train models for different tasks. For example, some clients may need to train models for image classification, while some other clients may need to train models for semantic segmentation [10]. Communication heterogeneity means that the participants of FL are usually under different network environments, and communication bandwidth and budgets of some participants are limited [11]. Therefore, limited bandwidth and unstable connection can significantly reduce the efficiency of FL. Device heterogeneity means that the clients vary in device performance, including computing power, storage resources, energy supply, etc. Since typical FL algorithms usually require several epochs of local training, the clients may update local models inconsistently and thus lower the efficiency of the whole system [12].
@@
Research addressing heterogeneous FL is currently generating a lot of buzz, and we investigate some highly cited and recent surveys related to heterogeneity and privacy. While they all provide detailed descriptions of their respective topics of focus, compared to our work, they overlook some less researched but equally important issues and techniques, and most of them cannot sufficiently cover both heterogeneity and privacy. In this paper, we first provide a detailed categorization of several types of heterogeneity and discuss the heterogeneity challenges that may be encountered when applying FL in real-world environments as well as state-of-the-art solutions. Considering that privacy preservation is one of the core requirements of FL, we discuss the application of existing privacy-preserving techniques to FL heterogeneous scenarios, exploring the trade-off relationship between privacy and utility, and presenting the latest research progress in related areas. We list the differences between our survey and the above other surveys in Table I. Thus, the main contributions of this work are threefold:
@@
Types of Heterogeneity Comparison of Surveys Representative Methods
[17] [18] [19] [20] [12] [21] [22] [23] [24] Ours
Data Heterogeneity Fig. 1 Distribution Skew $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ [25], [26], [27], [28], [29], [30], [10], [31], [32], [33], [34], [35], [36], [35]
Label Skew $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ [37], [26], [38]
Feature Skew $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ [39], [40], [41], [42], [41], [43]
Quality Skew $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ [44], [45], [24]
Quantity Skew $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ [46]
Model Heterogeneity Fig. 5 Structure Heterogeneity $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ [9], [47], [48], [49], [50], [51], [41], [52]
Optimizer Heterogeneity $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ [53]
Task Heterogeneity Fig. 2 $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ [10], [54], [47]
Communication Heterogeneity Fig. 5 $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ [40], [55], [49], [56], [57], [58], [59], [48], [60], [61]
Device Heterogeneity Fig. 5 $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ [32], [56], [55], [62], [57], [63], [64], [65], [66]
Privacy Mechanism Fig. 6 Differential Privacy $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ [67], [68], [69], [70], [71]
Homomorphic encryption $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ [72], [73]
Model Watermark $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ [74], [75], [76]
Blockchain $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ [77], [78], [79]
@@
    We systematically categorize and analyze key components in heterogeneous FL, including data, model, task, communication, and device heterogeneity. For each type, we discuss in detail the algorithmic challenges it raises. Comparing to previous surveys, we further refine some types of heterogeneity. For example, we further divide skews happened in label space into distribution skew and label skew, which are similar but may have different impact on the effect of FL.
@@
    Considering that some techniques can be applied to different types of heterogeneity, we discuss the FL methods in depth based on their adopted techniques, and categorize existing state-of-the-art methods into data-level, model-level, and architecture-level methods, and emphasize the types of heterogeneity targeted by each method.
@@
    We summarize the current privacy-preserving strategies in heterogeneous FL and explore possible future research directions with a view to promoting a safer and more efficient development of the field. Specifically, we discuss some privacy-preserving techniques that can be easily affected by heterogeneity as well as the corresponding countermeasures, which significantly matters but are not often explored in previous survey works.
@@ II Classification and Challenges in Heterogeneous Federated Learning
In this section, we illustrate the typical FL process through the equation. FL is ML setup in which multiple clients collaboratively train a model while protecting data privacy [67]. In a classic FL framework [2], assuming that there are $K$ clients participating in the training, for $k$-th client $C_{k}$ has a private dataset $D_{k}=\left\{(x_{i}^{k},y_{i}^{k})\right\}_{i=1}^{N_{k}}$ that satisfies $|x^{k}|=N_{k}$ and $N=\sum_{k=1}^{K}N_{k}$. The local distribution of the data of client $k$ is denoted as $P_{k}(x,y)$, where $x$ and $y$ denote the features of the data samples and the labels. Specifically, FL can be formally defined as the following optimization problem:
@@
where $\theta$ is the global model parameter. $F_{k}(\theta)$ is the local loss function on client $k$. $p_{k}$ is the client weight, the value of which depends on the amount of data on the client and fulfills $p_{k}=\frac{N_{k}}{N}$ and $\sum_{k=1}^{K}p_{k}=1$. Therefore, the training process of FL can be represented as the following steps:
@@
    $\theta\leftarrow\sum_{k=1}^{K}p_{k}\theta_{k},\quad\textbf{or}\quad\theta\leftarrow\theta+\sum_{k=1}^{K}p_{k}\Delta\theta_{k}.$ (2)
@@
Currently, FL has become increasingly widespread in real-world applications, especially in scenarios that require data privacy protection. In the healthcare sector [3], hospitals and medical institutions hold vast amounts of patient data, but due to privacy regulations, this data cannot be directly shared. FL allows these institutions to collaboratively train diagnostic models, such as disease prediction and medical image recognition, without exchanging sensitive data. In the financial industry [4], banks and financial institutions face a large volume of sensitive customer data (e.g., transaction records, credit information), which cannot be shared to prevent data breaches. FL enables multiple banks or financial institutions to collaboratively train models for fraud detection, credit scoring, and other applications while ensuring customer privacy is protected. In the Internet of Things (IoT) domain [5], smartphones, wearable devices, and other IoT devices generate vast amounts of data. These devices are often widely distributed and face limitations in computational resources and bandwidth. FL allows these devices to process data locally and train models, reducing data transmission costs, while optimizing personalized recommendation systems or speech recognition models without uploading personal data. In the field of intelligent transportation [6], autonomous vehicles can share knowledge between vehicles through FL without exchanging sensitive driving data. Different vehicles can locally train models to improve algorithms for path planning, driving decision-making, and traffic prediction, thus enhancing the safety and performance of the system. Despite the widespread use of FL in various fields [1], real-world FL systems often face challenges related to heterogeneity [17]. These include how to effectively deal with differences in data distribution across clients, how to adapt to the computational and storage capabilities of different clients while maintaining model performance, how to balance knowledge sharing with task specificity in multitasking scenarios, how to optimize communication strategies to cope with varying bandwidth and latency conditions, and how to adapt to the energy and resource constraints of different devices.
@@ II-A Data Heterogeneity
Data heterogeneity refers to the inconsistency of data distribution among clients in FL, which does not obey the same sampling, i.e., the data are not independently and identically distributed (Non-IID) [7, 24]. To explore data heterogeneity, we categorize data heterogeneity from the perspective of client data distribution [80]. Specifically, we distinguish different categories of Non-IID data: distribution skew, label skew, feature skew, quality skew, and quantity skew, as shown in Fig. 1. Data heterogeneity may cause the local model to converge in different directions to reach the local optimum rather than the global optimum, which reduces the FL performance and leads to poor global convergence of the model [81].
@@ II-A1 Distribution Skew
In FL, data distribution skew refers to the skew in the distribution of data held by different clients, i.e., the distribution of features or joint distribution of data from different clients is inconsistent with the overall data distribution [82]. In this case, the data no longer obeys the assumption of independent identically distributed (IID). As shown in Fig. 1a, there is $P_{i}(x,y)\neq P_{j}(x,y)$ due to the difference in the data distribution of clients $i$ and $j$. A typical example is data obeying a Dirichlet distribution [25] In the context of FL, the Dirichlet distribution can be used to model differences in the distribution of data labels on different clients. Where the parameter $\beta$ determines the degree of distributional difference of the data. As shown in the scatterplot in Fig. 1a, which illustrates the visualization of the difference in the distribution of data across 20 clients for different values of $\beta$.
@@ II-A2 Label Skew
Data label skew is a common Non-IID case, which refers to the skew in the distribution of data labels held by different clients. This label distribution skew means that even if the feature distributions are shared, the label distributions of different clients may be different [83], i.e., for clients $i$ and $j$, their label distributions $P_{i}(y)\neq P_{j}(y)$. However, for each label, the feature distributions of the data samples are similar regardless of which client they belong to, i.e., $P_{i}(x|y)=P_{j}(x|y)$. For example, in the data label skewing scenario shown in Fig. 1b, different users contain data from the set of handwritten digits, but each client may mainly contain a certain number of digits. The scatterplot shows a visualization of label skew for 20 clients. This data label skew can lead to a large difference in the gradient of the model uploaded back from different clients during federated training, thus affecting the convergence and generalization performance of the model.